One number is equal to the square of another. Find the numbers if both are positive and their sum is
132

Answer by Boreal(15235)

x=one number
x=y^2
x=y=132
y^2+y=132 by substitution
y^2+y-132=0
(y+12)(y-11)=0
only positive root is y=11
the numbers are 11 and 121; 121 is the square of 11.